OpenStudy (loser66):

\[J=\int_0^1\sqrt{1-x^4}\] \[K=\int_0^1\sqrt{1+x^4}\] \[L=\int_0^1\sqrt{1-x^8}\] A) J

OpenStudy (theopenstudyowl):

#walframa

OpenStudy (loser66):

Thanks, but no.

OpenStudy (meowlover17):

lol @theopenstudyowl

OpenStudy (theopenstudyowl):

Meow...lol

OpenStudy (meowlover17):

=^-^=

OpenStudy (loser66):

Please, dont' mess my post

OpenStudy (meowlover17):

@welshfella

OpenStudy (loser66):

@ganeshie8 any shortest way to find the answer without graphing or calculating?

OpenStudy (nincompoop):

visualise it

OpenStudy (loser66):

graphing?

ganeshie8 (ganeshie8):

Ahh graphing looks neat!

OpenStudy (nincompoop):

you don't want to "graph" it, so visualize

ganeshie8 (ganeshie8):

ohk i misread you don't want to graph/evaluate...

OpenStudy (loser66):

@nincompoop what is difference between graph and visualize?

ganeshie8 (ganeshie8):

visualize = graphing in ur head

OpenStudy (nincompoop):

GRAPH WOULD INVOLVE YOU ACTUALLY GRAPHING IT, VISUALIZE THAT INVOLVES JUST MENTAL

OpenStudy (anonymous):

Wouldn't \(\sqrt{1-x^{8}}\) have steep climbs and descents than \(\sqrt{1-x^{4}}\) ? I would expect because of the increased steepness in \(\sqrt{1-x^{8}}\) that it's integral would be of greater value. \(\sqrt{1+x^{4}}\) Would never tend towards the x-axis. It should be very similar to \(\sqrt{1-x^{4}}\) But increasing as x increases. So trying to logic it out, I would say J < L < K

OpenStudy (nincompoop):

THERE YOU GO!

ganeshie8 (ganeshie8):

For \(x\in (0,1)\) do we have \(x^m \lt x^n\) if \(m \gt n\) ?

ganeshie8 (ganeshie8):

then apply concentrationalizing logic

OpenStudy (loser66):

Got it. Thanks you all. :)

OpenStudy (anonymous):

You're welcome :)

ganeshie8 (ganeshie8):

looks the question boils down to simply putting the given integrands in ascending order

OpenStudy (loser66):

When graphing, J >1, and others <1, hence A is a correct answer. But I would like to shorten the time solving it. :)

ganeshie8 (ganeshie8):

\[x^4 \gt x^8 \implies -x^4 \lt -x^8 \implies 1-x^4\lt 1-x^8 \implies \sqrt{ 1-x^4}\lt \sqrt{ 1-x^8}\]

OpenStudy (loser66):

yyyyyyyyyyyyyyyes!! that 's good.

OpenStudy (loser66):

just from 0 to 1, that's true, right?

ganeshie8 (ganeshie8):

oh forgot to put that, yes all that is valid only for \(x\in (0,1)\)

OpenStudy (loser66):

Thanks a lot.

ganeshie8 (ganeshie8):

np sometimes it is hard to see \(x^4 \lt x\) for \(x\in (0,1)\)

ganeshie8 (ganeshie8):

if we let \(x = 0.9\), we get \(x^4 = 0.9*0.9*0.9*x\) which is clearly less than \(x\)

ganeshie8 (ganeshie8):

it is less than \(x\) because multiplying anything by a proper fraction always produces a smaller number (0.9 = 9/10 is a proper fraction)

OpenStudy (loser66):

yeah, and we know that 1+ whatever >1, hence K>1 , others <1, good interpretation. :)

ganeshie8 (ganeshie8):

yeah that was easy :)